# Reflexive Guidance: Improving OoDD in Vision-Language Models via Self-Guided Image-Adaptive Concept Generation

### v1

@@ C.1 Suggested Class Post-processing
Table C.2.4: ReGuide results on inputs that produce valid responses across three models on $5\%$ subset of the ImageNet200 benchmark. The AURC values are multiplied by $10^{3}$, and the other values are percentages. ‘Valid’ indicates the ratio of valid responses out of a total of $4,170$ image-prompt queries. Bold denotes the best performance among the results from each model.
Models ID Near-OoD Far-OoD All OoD
ImageNet200 NINCO SSB-Hard iNaturalist Textures Openimage-O
InternVL2-26B + ReGuide 65.46 (2,730) 95.1 10.67 36.65 23.87 / 86.42 15.02 / 90.92 1.54 / 98.3 3.21 / 97.80 2.11 / 97.99 10.22 / 93.6
InternVL2-76B + ReGuide 65.46 (2,730) 93.47 1.33 7.76 56.13 / 92.46 66.8 / 96.13 63.58 / 96.13 58.85 / 94.23 64.88 / 94.97 64.71 / 92.01
GPT-4o + ReGuide 65.46 (2,730) 91.43 1.33 10.13 98.06 / 96.16 99.31 / 96.6 99.69 / 96.6 98.93 / 95.24 98.27 / 96.29 99.09 / 94.01
C.3 Analysis on OoD Scores with ReGuide
Fig. C.3.1 presents the ROC curves for InternVL2-26B and GPT-4o, both with and without ReGuide. As shown in Fig. C.3.1(d), GPT-4o+ReGuide becomes slightly overconfident compared to GPT-4o without ReGuide. At TPR values near $100.00\%$ and $0.0\%$, the ROC curve indicates cases where OoD inputs are assigned higher confidence scores than ID inputs. Similarly, ROC on InternVL2-26B+ReGuide (Fig. C.3.1(b)) follows a pattern similar to that of GPT-4o+ReGuide. However, the FPR for InternVL2-26B+ReGuide is lower than that of InternVL2-26B without ReGuide, as the latter shows more instances where ID inputs receive lower confidence scores than OoD inputs, as depicted in Fig. C.3.1(a).
Tab.  C.3.1 presents cases where the model exhibits overconfidence, specifically highlighting samples at the two extremes –$100.00\%$ and $0.0\%$– of the ROC curve. These cases of overconfidence can be broadly categorized into two main issues. The first issue is OpenOoD labeling noise. As shown in Tab.  C.3.1, there are cases where an object occupying a small portion of the image is labeled as the ground truth instead of the dominant object. For example, InternVL2-26B suggests flower-related classes, while the ground truth is a bee. In the case of label noise, employing ReGuide can result in assigning high confidence to incorrect classes, leading to overconfidence. The second issue relates to the inherent characteristics of LVLMs. The models tend to assign high confidence to the most probable class rather than distributing confidence across multiple plausible classes. For instance, although many plausible classes are suggested in the Near-OoD category, GPT-4o assigns high confidence to only one class. InternVL2-26B shows a similar tendency. This overconfidence in a single class, rather than a broader distribution of confidence across possible classes, highlights a fundamental limitation of LVLMs.
@@ C.3 Analysis on OoD Scores with ReGuide
Table C.3.1: Samples at the extremes of the ROC curves for the ImageNet200 benchmark. Bold indicates the predicted class and its confidence score before applying softmax. Classes with a confidence score of $0.0$ have been omitted.
@@
0.00% [Uncaptioned image] ball $\bullet$ Dataset: SSB Hard
@@
Fig. C.3.2(a) and C.3.2(b) depict the OoD score distributions for InternVL2-26B and GPT-4o, respectively. As shown in Fig. C.3.2(a), with ReGuide, InternVL2-26B produces OoD scores ranging from $0.00$ and $100.00$, which is not observed without ReGuide. However, the effect of ReGuide on GPT-4o differs from that on InternVL2-26B. The OoD score distribution for GPT-4o becomes more biased with ReGuide, as illustrated in Fig. C.3.2(b). The number of responses with OoD scores between $0.00$ and $100.00$ decreases significantly. We infer that this outcome is due to the strong visual recognition ability of GPT-4o. As shown in Tab. 2, the suggestions from GPT-4o include classes that are highly similar to the given image.
@@
For a more detailed analysis of the ReGuide effect, we examine the OoD score distributions for ID and OoD inputs separately. Fig. C.3.3 displays the OoD score distributions for ID and OoD inputs on InternVL2-26B and GPT-4o. A common observation between InternVL2-26B and GPT-4o is the shift in the OoD score distribution of OoD inputs. As shown in Fig. C.3.3(b) and C.3.3(d), ReGuide increases the number of OoD inputs assigned an OoD score of 00, indicating that the models successfully classify these inputs into one of the suggested OoD classes, thereby enhancing OoD detectability. This trend is also observed, although to a lesser extent, for ID inputs in both models, as presented in Fig. C.3.3(a) and C.3.3(c). This reveals the underlying cause of the overconfidence behaviors observed in Fig. C.3.1(b) and Fig. C.3.1(d). The increasing number of ID inputs with low OoD scores leads to a misalignment between the OoD scores of ID and OoD inputs. We infer that this is due to some near-OoD class suggestions being overly similar or nearly synonymous with the ID classes, causing ID inputs to be misclassified into the suggested OoD classes. Despite this, ID accuracy with ReGuide improves compared to the models without ReGuide, as shown in Tab. 4.
@@ C.4 Comparison of OoD Score Functions
When incorporating available auxiliary OoD classes, one effective scoring strategy is to include confidence estimates for these auxiliary OoD classes. For example, NegLabel (Jiang et al., 2024) uses a scoring function that calculates the ratio of the sum of ID softmax values to the sum of both ID and OoD softmax values, while EOE (Cao et al., 2024) utilizes the difference between the maximum ID softmax and the maximum OoD softmax.
We compare three different scoring functions: 1) the maximum ID softmax, 2) the difference between the sum of ID softmax and OoD softmax, and 3) the difference between the maximum ID softmax and maximum OoD softmax. Each scoring function is represented in order as follows:
$\mathcal{S}_{\max}(\mathbf{x},\mathcal{Y},\mathcal{A}_{\text{aux}})=\max_{i\in\mathcal{Y}}\frac{e^{s_{i}/\tau}}{\sum_{k\in\mathcal{Y}\cup\mathcal{A}_{\text{aux}}}{e^{s_{k}/\tau}}}$ (1)
@@
Tab. C.4.1 presents the results using these scoring functions. All results are based on a $5\%$ subset of the ImageNet200 benchmark using InternVL2-26B+ReGuide. No significant differences are observed among the scoring functions. LVLMs tend to assign high confidence to a single predicted class, resulting in similar results regardless of the different scoring designs. Therefore, all results reported in our paper are measured by $\mathcal{S}_{\text{max}}$.
Table C.4.1: Comparison results of different scoring functions on a $5\%$ subset of ImageNet200 benchmark using InternVL2-26B
@@
$\mathcal{S}_{\text{max}}$ 22.92 / 86.60 16.25 / 89.98 1.23 / 98.19 3.81 / 97.15 2.23 / 97.67 10.79 / 93.01
$\mathcal{S}_{\text{sumsub}}$ 23.44 / 86.70 16.25 / 90.23 10.84 / 97.75 3.33 / 97.21 2.55 / 97.82 12.17 / 93.12
$\mathcal{S}_{\text{maxsub}}$ 23.44 / 86.73 16.32 / 90.23 1.23 / 98.29 3.81 / 97.22 2.23 / 97.82 10.86 / 93.20
@@ C.5 Prompt Format
We provide the full prompts used for experiments in Sec. 4.1. In the prompt, yellow, green, red, and blue boxes indicate a task description, an explanation of the rejection class, guidelines, and examples for the response format, respectively.
@@ C.5.1 Prompt for Class Suggestion of GPT-text
Fig. C.5.1 illustrates the prompt used for class suggestion of GPT-text. The guidelines consist of two key instructions: first, to avoid suggesting classes that are too similar to the image, and second, to avoid suggesting classes that are identical to the ID class. The red text indicates that the contents vary depending on the selected benchmark.
@@ C.5.2 Prompt for ReGuide Stage 1
Fig. C.5.2 illustrates the prompt used for class suggestion of ReGuide Stage 1. Similar to the original prompt used for evaluating OoDD and the prompt used for class suggestion in GPT-text, we provide guidelines to prevent uninformative class names, such as ID class names, overly similar classes, or simple repetitions of class names. Without an example, the response format of image-based suggestions tends to deviate more than that of text-based suggestions. Thus, we include the response format as an illustrative example. The red text indicates that the contents vary depending on the selected benchmark.
@@ C.5.3 Prompt for ReGuide Stage 2
Fig. C.5.3 illustrates the prompt used for ReGuide Stage 2. The red text indicates that the contents vary depending on the selected benchmark. The blue highlights the locations of the suggested OoD classes from Stage 1. The highlighted text in yellow indicates modifications made to incorporate these suggested classes. GPT-text employs the same prompt format for the classes suggested by GPT-4o.
@@
C.6 Limitations of ReGuide

### v2

@@ C.1 Suggested Class Post-processing
C.3 Analysis on OoD Score with ReGuide
To further investigate the FPR increase in larger models such as InternVL2-76B and GPT-4o in Tab. 4, we analyze the ROC curves for each model with and without ReGuide. Fig. C.3.1 shows the ROC curves for InternVL2-26B and GPT-4o. Both models exhibit improved ROC curves after applying ReGuide; however, the FPR declines more gradually for GPT-4o+ReGuide compared to InternVL2-26B+ReGuide as TPR decreases from $100\%$ as shown in Fig. C.3.1(b) and C.3.1(d). While the reduction in FPR for GPT-4o+ReGuide is less pronounced at high TPR thresholds (e.g., 100%–95%), it achieves significantly lower FPR as the TPR threshold approaches 90%.
@@ C.3 Analysis on OoD Score with ReGuide
To explore the causes of the FPR change observed across models in Fig. C.3.1, we analyze examples at the two extremes –$100.00\%$ and $0.0\%$– of the ROC curves, as well as the shifts in OoD score distributions induced by ReGuide. This analysis reveals that the increase in FPR@95%TPR for larger models such as InternVL2-76B and GPT-4o stems from their superior visual recognition ability, which leads to more confident predictions with highly suitable class suggestions.
Tab. C.3.1 presents examples on the extremes of Fig. C.3.1(b) and C.3.1(d). This highlights that labeling issues arise from insufficiently thorough annotations. First, a small subset of datasets exhibits a conceptual mismatch between the objects in images and their ground-truth labels. For instance, the grand piano image demonstrates a misprediction caused by such labeling issues, as the object is more accurately described as a grand piano mechanism, a class suggested by GPT-4o, rather than the ground-truth label grand piano. Another case involves images with ground-truth labels corresponding to objects that occupy only a small portion of the image, as shown in the bee and ball examples. In the bee image, the dominant object is a plant resembling a cactus, yet the ground-truth label is bee. For such images, LVLMs often suggest classes based on the dominant objects and classify them accordingly.
Table C.3.1: Examples at the extremes of the ROC curves for the ImageNet200 benchmark. Bold indicates the predicted class and its confidence score before applying the softmax function. Classes with a confidence score of $0.00$ have been omitted.
@@
0.00% [Uncaptioned image] ball $\bullet$ Dataset: SSB-Hard
@@
The suggested classes provided by GPT-4o for label-mismatched inputs are significantly fine-grained and contextually appropriate. Tab. C.3.2 presents the classes suggested by GPT-4o for images with ambiguous or inaccurate labels. For example, GPT-4o suggests classes such as model ape, gorilla figurine, orangutan statue, and monkey figurine as near-OoD classes. Based on the shape and texture of the object in the image, these suggestions are highly relevant and more suitable classifications. Similarly, for the cockshell image, GPT-4o suggests near-OoD classes such as motorboat, sailboat, rowboat, and kayak, which are fine-grained categories closely related to the depicted object. As shown in Tab. C.3.2, GPT-4o’s strong image recognition capabilities allow it to provide fine-grained, contextually appropriate classifications for images with ambiguous or inaccurate labels.
Table C.3.2: GPT-4o+ReGuide examples for images with ambiguous or inaccurate labels from the ImageNet200 benchmark. Bold indicates the predicted class and its confidence score before applying the softmax function. Classes with a confidence score of $0.00$ have been omitted.
ID inputs classified into OoD classes OoD inputs classified into ID classes
[Uncaptioned image] canoe $\bullet$ Dataset: ImageNet200
$\bullet$ Near-OoD: rowing boat, paddle, life jacket, riverbank, rapids, raft, white water rafting, paddling, inflatable boat, kayaking gear, river rafting, inflatables, oar, water sports, kayak, watercraft, stream
$\bullet$ Far-OoD: umbrella, outdoor adventure, bicycle, wristwatch, electric fan, soccer cleat, tennis racket, helmet, refrigerator, telescope, skateboard, basketball hoop, vacuum cleaner, washing machine, piano, laptop, microwave, desk chair, bookcase, coffee maker, alarm clock
$\bullet$ Prediction: kayaking gear
[…]
[Uncaptioned image] cockshell $\bullet$ Dataset: SSH-Hard
$\bullet$ Near-OoD: ferry, catamaran, life jacket, motorboat, sailboat, dinghy, dock, raft, rowboat, houseboat, life raft, trawler, paddleboat, gondola, skiff, marina, yacht, canoe paddle, pontoon boat, kayak
$\bullet$ Far-OoD: umbrella, bicycle, treadmill, electric fan, stethoscope, refrigerator, recliner, telescope, skateboard, typewriter, basketball hoop, vacuum cleaner, washing machine, sewing machine, suitcase, microwave, toaster, briefcase, alarm clock
$\bullet$ Prediction: canoe
[…]
[Uncaptioned image] chimpanzee $\bullet$ Dataset: ImageNet200
$\bullet$ Near-OoD: model ape, toy ape, gorilla figurine, primate figure, collectible primate, action figure chimp, novelty chimp, monkey figurine, toy baboon, orangutan statue, anthropoid toy, plastic monkey, chimpanzee toy, plush monkey, stuffed chimpanzee, baboon replica, simian statue, caricature gorilla, ape sculpture, monkey model
$\bullet$ Far-OoD: umbrella, bicycle, headphones, lamp, book, shoes, clock, refrigerator, pencil, guitar, computer, notebook, car, chair, printer, hat, toaster, sunglasses
$\bullet$ Prediction: monkey figurine
[…]
[Uncaptioned image] boar $\bullet$ Dataset: SSB-Hard
$\bullet$ Near-OoD: hog, boar, peccary, berkshire pig, potbellied pig, saddleback pig, warthog, hampshire pig, pietrain pig, wild boar, sow, yorkshire pig, javelina, meishan pig, tamworth pig, duroc pig, mangalitsa pig, kune kune, vietnamese pig, large white pig
$\bullet$ Far-OoD: umbrella, wristwatch, bicycle, electric fan, soccer cleat, refrigerator, telescope, skateboard, bookcase, basketball hoop, vacuum cleaner, washing machine, piano, laptop, microwave, tennis racket, desk chair, coffee maker, alarm clock
$\bullet$ Prediction: pig
[…]
[Uncaptioned image] wheelbarrow $\bullet$ Dataset: ImageNet200
$\bullet$ Near-OoD: caster, cart, wheel rim, unicycle, wagon, trolly wheel, bicycle wheel, pushcart, hand truck, scooter, garden barrow, dolly, wheel, yard cart, stroller, tricycle, garden trolley, motorcycle wheel, tire
$\bullet$ Far-OoD: umbrella, wristwatch, electric fan, paintbrush, soccer cleat, tennis racket, electric drill, refrigerator, telescope, skateboard, basketball hoop, vacuum cleaner, washing machine, piano, laptop, microwave, desk chair, bookcase, coffee maker, alarm clock
$\bullet$ Prediction: wheelbarrow
[…]
[Uncaptioned image] peacock $\bullet$ Dataset: SSB-Hard
$\bullet$ Near-OoD: damselfly, moth, wasp, termite, cricket, bumblebee, flea, aphid, hornet, lacewing, stink bug, praying mantis, beetle, cicada, firefly, katydid
$\bullet$ Far-OoD: umbrella, wristwatch, bicycle, electric fan, soccer cleat, tennis racket, refrigerator, telescope, skateboard, basketball hoop, vacuum cleaner, washing machine, piano, coffee maker, laptop, microwave, desk chair, bookcase, alarm clock
$\bullet$ Prediction: monarch butterfly
[…]
@@
Fig. C.3.2(a) and C.3.2(b) depict the OoD score distributions for InternVL2-26B and GPT-4o, respectively. As shown in Fig. C.3.2(a), with ReGuide, InternVL2-26B produces OoD scores ranging between $0.00$ and $100.00$, which is not observed without ReGuide. However, the effect of ReGuide on GPT-4o differs from that on InternVL2-26B. The OoD score distribution for GPT-4o becomes more biased with ReGuide, as illustrated in Fig. C.3.2(b). The number of responses with OoD scores between $0.00$ and $100.00$ decrease significantly. We infer that this outcome is due to the strong visual recognition ability of GPT-4o. As shown in Tab. 2, the suggestions of GPT-4o include classes that are highly similar to the given image.
@@
For more detailed analysis of the ReGuide effect, we examine the OoD score distributions for ID and OoD inputs separately. Fig. C.3.3 displays OoD score distributions for ID and OoD inputs on InternVL2-26B and GPT-4o.
@@
A common observation regarding ReGuide for InternVL2-26B and GPT-4o is the shift in the OoD score distribution of OoD inputs. As shown in Fig. C.3.3(b) and C.3.3(d), ReGuide increases the number of OoD inputs with an OoD score of $0.0$ while reducing those with a score of $100$. This shift demonstrates that the models effectively classify OoD inputs into one of the suggested OoD classes, thereby improving OoD detectability and confirming that ReGuide functions as intended.
The increase in inputs with an OoD score of $0.0$ is also observed for ID inputs. As previously discussed in Tab. C.3.1, this is caused by ID images with ambiguous or inaccurate labels. For such inputs, GPT-4o+ReGuide effectively classifies them into a single class with higher confidence, as the suggested classes often include one closely aligned with the ground-truth label. This explanation is further supported by the improvement in FPR@90%TPR, where the influence of this small subset of images is reduced, as shown in Tab. 4.
@@ C.4 Comparison of OoD Score Functions
When incorporating available auxiliary OoD classes, one of effective scoring strategies is to include confidence estimates for these auxiliary OoD classes. For example, NegLabel (Jiang et al., 2024) uses the scoring function as the ratio of the sum of ID softmax to the sum of both ID and OoD softmax, and EOE (Cao et al., 2024) utilizes the difference between max ID softmax and max OoD softmax.
We compare three different scoring functions: 1) the maximum ID softmax, 2) the difference between the sum of ID softmax and OoD softmax, and 3) the difference between the maximum ID softmax and maximum OoD softmax. Each of the scoring functions is represented, in order, as follows:
$\vspace{-10pt}\mathcal{S}_{\max}(\mathbf{x},\mathcal{Y},\mathcal{A}_{\text{aux}})=\max_{i\in\mathcal{Y}}\frac{e^{s_{i}/\tau}}{\sum_{k\in\mathcal{Y}\cup\mathcal{A}_{\text{aux}}}{e^{s_{k}/\tau}}}$ (1)
@@
Tab. C.4.1 presents the results with these scoring functions. All results are on a $5\%$ subset of ImageNet200 benchmark using InternVL2-26B+ReGuide. No significant differences are observed across the scoring functions. LVLMs tend to assign high confidence to a single predicted class, resulting in similar results regardless of these different scoring designs. Therefore, all results reported in our paper are measured by $\mathcal{S}_{\text{max}}$.
Table C.4.1: Comparison results between different scoring functions on a $5\%$ subset of ImageNet200 benchmark using InternVL2-26B
@@
$\mathcal{S}_{\text{max}}$ 22.89 / 86.53 15.21 / 90.41 1.39 / 98.02 3.93 / 97.05 2.04 / 97.68 10.27 / 93.19
$\mathcal{S}_{\text{sumsub}}$ 23.88 / 86.46 15.21 / 90.58 10.65 / 97.45 3.49 / 96.90 2.33 / 97.69 11.59 / 93.19
$\mathcal{S}_{\text{maxsub}}$ 23.38 / 86.53 15.27 / 90.59 1.39 / 97.45 3.93 / 96.90 2.04 / 97.69 10.33 / 93.26
@@ C.5 Prompt Format
We provide the full prompts used for experiments in Sec. 4.1. In the prompt, yellow, green, red, blue boxes indicate a task description, an explanation of rejection class, guidelines, and examples for response format, respectively.
@@ C.5.1 Prompt for Class Suggestion of GPT-text
Fig. C.5.1 illustrates the prompt used for class suggestion of GPT-text. The guidelines consist of two key instructions: first, to avoid suggesting classes that are too similar to the image, and second, to avoid suggesting classes that are identical to the ID class. The red text indicates that the contents vary depending on the choice of benchmark.
@@ C.5.2 Prompt for ReGuide Stage 1
Fig. C.5.2 illustrates the prompt used for class suggestion of ReGuide Stage 1. Similar to the original prompt used for evaluating OoDD and the prompt used for class suggestion on GPT-text, we provide a few guidelines to prevent receiving uninformative class names, such as ID class names, overly similar class names or just a repetition of a class name. Without an example, the response format of image-based suggestions deviates more compared to that of text-based suggestions. Thus, we include the response format as an illustrative example. The red text indicates that the contents vary depending on the choice of benchmark.
@@ C.5.3 Prompt for ReGuide Stage 2
Fig. C.5.3 illustrates the prompt used for ReGuide Stage 2. The red text indicates that the contents vary depending on the choice of benchmark. The blue indicates the locations of the suggested OoD classes from Stage 1. The highlighted text in yellow indicates modifications made to reflect these suggested classes. GPT-text employs the same prompt format for GPT-4o’s suggested classes.
@@
C.6 Inference cost analysis
While ReGuide, as a two-stage approach, incurs higher computational costs compared to single-turn generation (Section 3), the inference cost does not double. This is because caching image representations eliminates the need to process the same image twice, and the shorter prompt in Stage 1 (class suggestion) reduces the token count. For open-source models, we analyze the inference cost from a computational complexity perspective, whereas for proprietary models, the analysis focuses on API usage costs.
We estimate the inference time $FLOPs$ following Li et al. (2024b). This standard practical estimate is based on a simplified scaling law, assuming that the inference cost of the LLM scales linearly with both the number of tokens processed and the model size. The inference time $FLOPs$ is estimated as:
$FLOPs=\mathcal{O}(N\times T)$ (4)
where $N$ is the number of parameters in the LLM and $T$ denotes the total number of tokens involved in the inference process. Specifically, $T$ is the sum of $Q$, $V$, and $G$, which correspond to the number of text input tokens, image tokens, and generated tokens, respectively. For example, for InternVL2-76B, the prompt depicted in Fig. B.6.1 with ImageNet200 classes includes 5,350 input tokens. The number of generated tokens is 1,936, assuming the output follows the provided example structures, and the number of image tokens is 256. Thus, the baseline $FLOPs$ is estimated as $7\times 10^{10}\times(256+5,350+1,936)$.
Similarly, for ReGuide, $FLOPs$ is estimated as:
$FLOPs=\mathcal{O}(N\times(V+Q_{Stage1}+G_{Stage1}+Q_{Stage2}+G_{Stage2}))$ (5)
where $Q_{Stage1},G_{Stage1}$ and $Q_{Stage2},G_{Stage2}$ denote the number of text input and generated tokens for Stage 1 and Stage 2, respectively. Although ReGuide involves two stages, the input image is processed only once as the same image is reused across both stages and can be cached. Note that the inference cost of the vision encoder is excluded, as it remains consistent for both the baseline and ReGuide within the same VLM.
Tab. C.6.1 presents $FLOPs$ for the baseline and ReGuide on InternVL2 models. These values are estimated on one input pair (image-prompt) of the ImageNet200 benchmark, assuming that the generated outputs follow one of the provided example structures.
Table C.6.1: Estimated computational costs for InternVL2 models (in units of $10^{13}$ FLOPs)
Models Baseline +ReGuide
InternVL2-26B 15.086 17.052
InternVL2-76B 52.794 60.144
    InternVL2-InternLM2-Chat-26B, InternVL2-LLaMA3-76B
ReGuide requires only 1.14$\times$ the computational cost compared to the baseline. It is because 1) the prompt used in Stage 1 is short and simple (e.g., $917$ tokens for Stage 1), and 2) the image tokens are processed only once. Consequently, the increase in computational cost arises solely from the processing of the Stage 1 prompt. Additional optimization techniques for multi-turn inference could further reduce the computational cost gap between the baseline and ReGuide.
Tab. C.6.2 shows API costs for proprietary models under the baseline and ReGuide experiments on the 5% ImageNet200 benchmark. From the perspective of API costs, ReGuide exhibited a higher proportional increase compared to computational costs above, with a 1.47$\times$ increase relative to the baseline. Since we processed each stage separately, the API cost reflects a scenario where image token caching is not applied. Additionally, we did not employ multi-turn inference, as it provided no advantages in inference time due to the absence of batch inference support. Implementing multi-turn batch inference could significantly improve efficiency and reduce costs.
Table C.6.2: API costs (USD) for proprietary models, calculated per $1$ million input and output tokens
Models Input token Output token Baseline ReGuide
GPT-4o 1.25 5 73 107
Claude 3.5 Sonnet 3 15 228 -
Gemini Pro 1.5 3.5 10.5 193 -
    GPT-4o (2024-08-06)
One possible approach to reduce the inference cost gap between the baseline and ReGuide is to integrate the two-stage prompts into a single prompt while maintaining ReGuide’s process of providing suggested classes. This adjustment would allow ReGuide to maintain its effectiveness with a single forward pass. Fundamentally, it is essential to develop LVLMs to overcome their tendency to limit responses to the options explicitly provided in the user-defined prompt. Such advancements would enable LVLMs to explore and utilize information beyond the immediate constraints of the prompt, establishing robust models capable of generating safe and reliable responses to a wider range of inputs.
C.7 Limitations of ReGuide